OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4) A. -2 B. 1 /2 C. -1/2 D. 2

OpenStudy (anonymous):

it is easly slope (m) = y2-y1/(x2-x1) hence putting this values we get m=(4-2)/(7-6) =2/1=2 hence answer is 2

OpenStudy (anonymous):

thank you both very much

OpenStudy (anonymous):

welcome!!!